Suppose you have a bag containing the following
32 red beads
18 white beads
12 black beads
24 yellow beads
8 pink beads..
Suppose you draw 2 beads out of the bag with replacement. Find the probability of drawing a red bead and then a black bead..

Answer by jim_thompson5910(35256)

There are 32+18+12+24+8 = 94 beads total

P(red then black) = P(red) * P(black)

P(red then black) = (32/94)*(12/94)

P(red then black) = (32*12)/(94*94)

P(red then black) = (384)/(8836)

P(red then black) = 96/2209

P(red then black) = 0.04345857854232 (note: this is approximate)